OpenStudy (schoolgeek15):

The sum of three consecutive numbers is 72. What is the largest of these numbers? 23 25 26 28

Vocaloid (vocaloid):

hint: let n = the first of the three numbers let n + 1 = the second number let n + 2 = the third number n + n + 1 + n + 2 = 72 solve for n

OpenStudy (schoolgeek15):

Would it be 24? That's what I got but its not and answer choice @Vocaloid

Vocaloid (vocaloid):

3n + 3 = 72 3n = 69 n = ?

OpenStudy (schoolgeek15):

25?

Vocaloid (vocaloid):

what is 69 divided by 3?

OpenStudy (schoolgeek15):

23

Vocaloid (vocaloid):

good, but we're not done yet

OpenStudy (schoolgeek15):

Oh!

Vocaloid (vocaloid):

the first number is 23 which means the second number is 24 and the last number is 25 so 25 is our answer

OpenStudy (schoolgeek15):

Oh! I see thank you very much!